OpenStudy (anonymous):

Complete the equation of the graphed linear function. Write the slope in decimal form. y = x + The graph of the function goes through point negative 5, 0 and point 0, 4.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@KyanTheDoodle

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@roz_mar23

OpenStudy (anonymous):

So you need the equation for Y=Mx+B?

OpenStudy (anonymous):

yes i need help for someone to walk me though it

OpenStudy (anonymous):

ok ill walk you through

OpenStudy (anonymous):

great

OpenStudy (anonymous):

i know we need to do y2-y1/x2-x1 its after that i get confused

OpenStudy (anonymous):

So to fine M or the slpoe you use the eqation \[\frac{ Y2-Y1 }{ X2-X1 }\] So that is \[\frac{ 4-0 }{ 0-5}\] that Equals -4/5

OpenStudy (anonymous):

So -4/5 is the slope

OpenStudy (anonymous):

thats the slope which we should put in decimal form

OpenStudy (anonymous):

as the question says lol

OpenStudy (anonymous):

-0.8 is the slope

OpenStudy (anonymous):

Yes correct

OpenStudy (anonymous):

so y=-0.8x+?

OpenStudy (anonymous):

Now the B is where the line crosses X

OpenStudy (anonymous):

thats were i get stuck

OpenStudy (anonymous):



OpenStudy (anonymous):

 Thats where the line crosses X so B equals 4